Incomplete manufacturing cost data for Horizon Company for 2017 are presented as follows for four different situations. Indicate the missing amount for each letter. Direct Materials Used Direct Labor Used Work in ProcesS Work in Process 12/31 Total Cost of Goods Manufactured Manufacturing Manufacturing Overhead $91,200 133,700 Costs (1) $122,700 $142,700 (a) $36,800 (b) $365,100 (c 203,200 452,400 268,100 293,500 43,000 475,600 85,900 105,000 64,500 83,700 73,200 80,400 46,300 Ch) 273,800 LINK TO TEXT VIDEO: APPLIED SKILLS Prepare a condensed cost of goods manufactured schedule for situation (1) for the year ended December 31, 2017 HORIZON COMPANY Cost of Goods Manufactured Schedule
Answer #1
Direct material used Direct labour used Manufacturing overhead Total manufacturing overhead Work in process 1/1 Work in process 12/31 Cost of goods manufactured
1 122700 142700 91200 356600(a) 36800 28300 (b) 365100
2 115500© 203200 133700 452400 19800(d) 43000 475600
3 85900 105000 77200( e) 268100 64500 83700 248900 (f)
4 73200 139900 (g) 80400 293500 46300 66000 (h) 273800
Horizon company
Cost of goods manufactured schedule
Work in process 1/1 36800
Direct material 122700
Direct labour used 142700
Manufacturing overhead 91200
Total manufacturing cost 356600
Total cost of WIP 393400
Less: WIP 31/12 28300
Cost of goods manufactured schedule 365100
(a) total manufacturing overhead = 122700+142700+91200 = 356600
(b) WIP 12/31 = 365100-356600-36800 = 28300
( c) Direct material used = 452400-133700-203200 = 115500
(d) WIP 1/1 = 475600-452400-43000 = 19800
(e ) Manufacturing overhead = 268100-85900-105000 = 77200
( f ) Cost of goods manufactured = 268100+64500-83700 = 248900
(g) Direct labour used = 293500-80400-73200 = 139900
(h) WIP 31/12 = 273800-293500-43600 = 66000